OpenStudy (anonymous):

A square is inscribed inside a circle. The square's vertices have coordinates R(0, 4), S(6, 2), T(4, -4), and U(-2, -2). Find the equation of the circle. (x - ? )^2 + y^2 = ?

hero (hero):

I'm not even going to attempt to help you since I've already failed twice

OpenStudy (anonymous):

(x-2)^2=y^2=sqrt80

OpenStudy (anonymous):

i think the sum is ronq

OpenStudy (mertsj):